# TREASUREHUNT: Transients and Variability Discovered with HST in the JWST North Ecliptic Pole Time Domain Field

### v2

@@ Abstract
The JWST North Ecliptic Pole (NEP) Time Domain Field (TDF) is a 14′ diameter field optimized for multi-wavelength time-domain science with JWST. It has been observed across the electromagnetic spectrum both from the ground and from space, including with the Hubble Space Telescope (HST). As part of HST observations over 3 cycles (the “TREASUREHUNT” program), deep images were obtained with ACS/WFC in F435W and F606W that cover almost the entire JWST NEP TDF. Many of the individual pointings of these programs partially overlap, allowing an initial assessment of the potential of this field for time-domain science with HST and JWST. The cumulative area of overlapping pointings is 88 arcmin, with time intervals between individual epochs that range between 1 day and 4 years. To a depth of 29.5 mag (F606W), we present the discovery of 12 transients and 190 variable candidates. For the variable candidates, we demonstrate that Gaussian statistics are applicable, and estimate that 80 are false positives. The majority of the transients will be supernovae, although at least two are likely quasars. Most variable candidates are AGN, where we find 0.42% of the general 6 field galaxy population to vary at the level. Based on a 5-year timeframe, this translates into a random supernova areal density of up to 0.07 transients per arcmin (245 deg) per epoch, and a variable AGN areal density of 1.25 variables per arcmin (4500 deg) to these depths.
@@ 1 Introduction
Jansen & Windhorst (2018) selected a particularly suitable area within JWST’s northern CVZ, the JWST North Ecliptic Pole (NEP) Time Domain Field (TDF), to enable the exploration of time-varying phenomena with JWST at high redshifts, but also within the halo of our own Galaxy, and in the extreme outer Solar System at high ecliptic latitudes. The JWST NEP TDF is a 14′ diameter field centered on (RA, Dec) = (17:22:47.896, +65:49:21.54), that was carefully chosen to minimize foregrounds, to be devoid of stars brighter than 15.5 mag, and maximize the observing efficiency for time-domain science with JWST. Its location near the NEP allows for follow-up of transients and variable sources at any cadence and at any time of the year. In addition, the NEP suffers minimal Zodiacal foregrounds, with minimal variation in the course of a year, naturally allowing for more sensitive observations per unit time. While the ecliptic latitude of the JWST NEP TDF is very high ( 86.2), it is located at intermediate Galactic latitude ( 33.6), providing a clear but relatively long sight-line through the halo of our Galaxy. This makes the field useful for deep Galactic time-domain science with JWST as well.
This field serves as an important testbed for advancing our understanding of the time-varying universe at fainter limits than can be reached from the ground at optical and near-IR wavelenghts, where transients and variability can hold cosmological significance (e.g., dark energy, cosmic star formation rate, and the evolution of supermassive blackholes). Type Ia SNe are often used as standard candles to provide critical constraints on the Hubble constant, the mass density of the universe, the cosmological constant, the deceleration parameter, and the age of the Universe (e.g., Riess et al., 1998). Core Collapse (CC) SNe are equally important, as they not only provide most of the heavy elements in our universe (e.g., Matteucci & Greggio, 1986), but also should trace the rate of instantaneous massive star formation. CC SNe rates and their connection to cosmic star formation rates remains poorly understood, as there seems to be a significant mismatch between observed CC SNe rates and what is expected (see, e.g., Cappellaro, 2014).
@@
While previous HST variability studies like Sarajedini et al. (2003) (Hubble Deep Field North) and Cohen et al. (2006) (Hubble Ultra Deep Field) achieved impressive depths (29.0 and 30.5 mag, respectively), they are limited to approximately 6.5 arcmin. Villforth et al. (2010), on the other hand, leveraged the expansive area (500 arcmin) of the Great Observatories Origins Deep Survey (GOODS), but to a relatively shallow depth (26.0 mag). At the same exquisite angular resolution (full width half max 009), TREASUREHUNT observations provide a unique combination of depth ( 29.5 mag) and area (88 arcmin) for a variability analysis.
In this work, we provide the first time-domain study of the JWST NEP TDF at visible wavelengths, presenting 12 transients (supernova candidates) and 100 variable candidates (AGN candidates) identified using TREASUREHUNT data. In § 2, we briefly summarize the TREASUREHUNT HST data, and data processing specific to the detection of transients and variable objects. § 3 explains our methods for identifying transients and variability in the NEP TDF. In § 4.1 and § 4.2 we showcase our supernova candidate detections and our variable AGN candidates, respectively. All magnitudes are in AB units222Defined as mag, where the flux density is in Jy. (Oke & Gunn, 1983). Where relevant, we adopt a flat CDM cosmology with km s Mpc, , and (Planck Collaboration et al., 2016, 2020).
@@ 2.1 HST Observations and Data Processing
Observations from HST GO 15278 were taken between 2017 October 1 and 2019 February 9, and those from GO 16252+16793 (TREASUREHUNT333We will loosely refer to the observations and data of the GO 15278 program as also part of TREASUREHUNT hereafter, as the intent and survey strategy were identical.) between 2020 September 25 and 2022 October 31. The former program targeted the central ( 5′) portion of the JWST NEP TDF, and the latter an outer annulus to 78. Together, these programs provide near-contiguous coverage with the Wide Field Camera 3 (WFC3/UVIS) in the F275W filter ( 0.272 µm) and with the Advanced Camera for Surveys (ACS/WFC) in the F435W and F606W filters ( 0.433 and 0.592 µm), with a total area of 194 arcmin (ACS/WFC). The full HST coverage of the JWST NEP TDF is shown in Figure 1.
As described in more detail in Jansen et al. (2024a,b; in prep.), both programs used 4-orbit CVZ and pseudo-CVZ444True CVZ orbits imply Earth Limb angles 15.5, but uninterrupted observations can be implemented at smaller angles down to 7.6 away from the Earth limb when phased to execute when that limb is dark. We term these possible but non-standard opportunities “pseudo-CVZ”. visits to reach 2- limiting depths of 28.0, 28.6, and 29.5 mag in F275W, F435W and F606W, respectively. These limiting depths were determined using the SourceExtractor (Bertin & Arnouts, 1996) MAG_AUTO aperture, within areas with single-epoch coverage (lighter grey in Figure 1), and are the adopted nominal survey depths.
@@
The background-subtracted images were corrected for geometric distortion and aligned to the Gaia/DR3 (Gaia Collaboration et al., 2016, 2023) astrometric reference grid using the DrizzlePac tasks TweakReg and AstroDrizzle (see Hoffmann et al., 2021). The AstroDrizzle software also flags cosmic-rays, a crucial step when identifying faint variability, where cosmic ray induced signals could cause spurious variability if not properly removed. We expect, and tested (as discussed in more detail in Appendix A), that all cosmic ray and other image artifacts were largely removed through the AstroDrizzle step, as each single 4-orbit visit resulted in 8–9 exposures per filter that were stacked together.
All 8–9 exposures per visit were taken at the same pointing with small 0.2″ dither offsets in RA and Dec. For each visit, we drizzle-combined the individual post-processed, rectified and aligned exposures per filter (Jansen et al. 2024a,b; in prep.). For ACS/WFC F606W and F435W, this resulted in 24 rectified and aligned science images (and 24 associated weight maps), and for WFC3/UVIS F275W in 23 images and associated weight maps, each at a pixel scale of 0030 pix and centered on the nominal field center of the JWST NEP TDF. We also use the full mosaics of all drizzle-combined images per filter, generated with the same image dimensions and pixel scale, and aligned to Gaia/DR3 to within 0008 in RA and Dec.
The footprints of individual ACS/WFC (and to a lesser extent WFC3/UVIS) visits partially overlap one another, with a cumulative area of overlap of 88 arcmin. These overlapping regions are key to searching for transients and variability within this field. Figure 1 showcases the overlapping regions in progressively darker shades of grey for areas of 2-, 3- and 4-epoch overlap. The time interval between observations in these regions of overlap range between 1 day and 4 years.
We note that while the very small number of epochs and sparse time sampling can catch transients like SNe and variable sources like AGN, it cannot provide a light curve to trace their evolution. Further epochs would provide a more robust identification of variable sources, where we may characterize the scatter in the measured brightness (e.g., Pouliasis et al., 2019; Sokolovsky et al., 2017).
@@ 2.2 JWST Observations and Data Processing
JWST/NIRCam observations within the JWST NEP TDF were obtained between 2022 August 25 and 2023 May 30 as part of JWST GTO 2738 (PIs: R. Windhorst & H. Hammel) in 7 broadband filters (F090W through F444W) and 1 mediumband filter (F410M) as part of the PEARLS program (Windhorst et al., 2023). Figure 1 shows the JWST observations outlined in orange, highlighting the 4 epochs, corresponding to the four orthogonal “spokes”. Each spoke consists of two NIRCam pointings that partially overlap in the center. Each pointing includes a total exposure time of 3.5 hours across all 8 filters, or about 3000 sec per filter. Each spoke has an area of 137, totalling 547 of JWST observations.
The data were retrieved from MAST and post-processed by the PEARLS team using their custom pipeline to mitigate noise, wisps in the NIRCam/SW filters, mask snowball artifacts, and flatten the background across read-out amplifier boundaries. The individual post-processed images where rectified and aligned to Gaia/DR3 independent of the HST images, but using nearly identical methods. Full mosaics of the field were created for each filter with a 003 platescale. The 5 sigma point-source limit is typically between 28.0 and 29.1 mag, depending on the filter. For more details on the reduction, calibration and post-processing, we refer to Windhorst et al. (2023), Robotham et al. (2023), and Jansen et al. 2024c (in prep.). In § 4.3, we will use PEARLS aperture photometry of transient and varying objects discovered in areas with both HST and JWST coverage to fit SEDs.
@@ 2.3 X-Ray Observations and Data Processing
The JWST NEP TDF has also been observed extensively in X-rays with NuSTAR, XMM-Newton, and Chandra. NuSTAR has monitored the field since Cycle 5 (PID 5192, 6218, 8180, and 9267; PI: F. Civano). For this work we use the Cycle 5+6 data and catalog published in Zhao et al. (2021a, b, 2024). The 21 NuSTAR observations took place over 28 months (between 2019 September and 2022 January), while the three XMM-Newton observations (part of the NuSTAR cycle 6 program) were taken over 15 months (between 2020 October and 2022 January). The total NuSTAR exposure from Cycle 5+6 is 1.5 Ms, reaching a flux limit of 3.310 erg cm s in the 3–24 keV band. We note that this is the deepest NuSTAR field taken thus far. The total XMM-Newton exposure time is 62 ks, reaching a limit of 8.710 erg cm s in the 0.5–2 keV band.
We also use the first 1.3 Ms of Chandra observations of the JWST NEP TDF (PID 19900666, 20900658, 21900294 and 22900038; PI: W.P. Maksym) spanning cycles 19–23. Chandra observed the field 46 times between 2018 April 12 and 2022 September 25 (a span of 4.6 years), with exposures ranging from 9 to 59.9 ks. Beginning 2021 August 29, these occurred in 90 ks epochs every 3 months, typically broken into 2 or more individual visits (OBSIDs) spanning no more than 30 days per epoch. We reprocessed and reduced the data using ciao (Fruscione et al., 2006), corrected the cross-observations using source catalogs generated by wavdetect, and astrometrically registered them to Pan-STARRS (Chambers et al., 2016). We then used merge_obs to generate combined event files and exposure maps. The faintest 3 detection in the catalog is erg s cm in the full 0.5–7 keV band assuming galactic absorption and a power law.
@@ 3 Methods to Identify Transients and Variability
In this section, we detail our methods for identifying transients and variability in the 55 areas of overlap (shown as darker shades of grey in Figure 1) with a combined area of 88 arcmin.
@@ 3.2 Variability Search
Note. For each SourceExtractor parameter in the first column, the second an third columns list its value for identifying variable sources and for full aperture photometry, respectively. We used a circular aperture of 024 diameter for variability identification, specified using the PHOT_APERTURES parameter, and the FLUX_AUTO aperture for full aperture photometry. We set the GAIN value to the median exposure time of the individual exposures combined per visit.
@@ 3.2.3 Accounting for False Detections
In addition to false detections due to Gaussian statistics, image artifacts like cosmic rays, hot pixels, bias stripping, amplifier offsets, CTE trails, dust motes, optical ghosts, and cross-talk, may cause artificial variability to be detected. Of these, cosmic rays are the brightest and most prevalent, yet will not affect our results due to robust flagging and having 8 exposures (per pointing, filter and epoch) to stack using the AstroDrizzle software. We explore this in more detail in Appendix A. The other image artifacts are largely removed by the standard HST pipeline. Therefore, we do not expect and find no evidence for image artifacts to bias our results. Nonetheless, confirmation of which individual sources genuinely vary will require follow-up observations.
@@ 4.1 Transients in the JWST NEP Time Domain Field
In all, we detected 0.14 transients per arcmin of overlapping area, or 491 deg. Since two epochs are needed to detect a transient, the areal density per epoch is therefore 0.07 transients per arcmin (245 deg), or 0.77 per ACS/WFC footprint per epoch. For comparison, Dahlen et al. (2012) searched for SNe in the GOODS fields (using the ACS/F850LP filter), and identified 118 SNe in a total area of 1 deg to 26 mag. If we only consider the 4 transients with 26.0 mag, we find an areal density of 164 deg, consistent with Dahlen et al. (2012) when factoring in differences in detection filter, SNe colors and redshift distribution, and small number statistics. We also note that it is likely that not all our transients are SNe.

### v3

@@ Abstract
The JWST North Ecliptic Pole (NEP) Time Domain Field (TDF) is a 14′ diameter field optimized for multi-wavelength time-domain science with JWST. It has been observed across the electromagnetic spectrum both from the ground and from space, including with the Hubble Space Telescope (HST). As part of HST observations over 3 cycles (the “TREASUREHUNT” program), deep images were obtained with ACS/WFC in F435W and F606W that cover almost the entire JWST NEP TDF. Many of the individual pointings of these programs partially overlap, allowing an initial assessment of the potential of this field for time-domain science with HST and JWST. The cumulative area of overlapping pointings is 88 arcmin2, with time intervals between individual epochs that range between 1 day and 4 years. To a depth of 29.5 mag (F606W), we present the discovery of 12 transients and 190 variable candidates. For the variable candidates, we demonstrate that Gaussian statistics are applicable, and estimate that 80 are false positives. The majority of the transients will be supernovae, although at least two are likely quasars. Most variable candidates are AGN, where we find 0.42% of the general 6 field galaxy population to vary at the level. Based on a 5-year timeframe, this translates into a random supernova areal density of up to 0.07 transients per arcmin2 (245 deg-2) per epoch, and a variable AGN areal density of 1.25 variables per arcmin2 (4500 deg-2) to these depths.
@@ 1 Introduction
Jansen & Windhorst (2018) selected a particularly suitable area within JWST’s northern CVZ, the JWST North Ecliptic Pole (NEP) Time Domain Field (TDF), to enable the exploration of time-varying phenomena with JWST at high redshifts, but also within the halo of our own Galaxy, and in the extreme outer Solar System at high ecliptic latitudes. The JWST NEP TDF is a 14′ diameter field centered on (RA, Dec) = (17:22:47.896, +65:49:21.54), that was carefully chosen to minimize foregrounds, to be devoid of stars brighter than 15.5 mag, and maximize the observing efficiency for time-domain science with JWST. Its location near the NEP allows for follow-up of transients and variable sources at any cadence and at any time of the year. In addition, the NEP suffers minimal Zodiacal foregrounds, with minimal variation in the course of a year, naturally allowing for more sensitive observations per unit time. While the ecliptic latitude of the JWST NEP TDF is very high ( 86.2∘), it is located at intermediate Galactic latitude ( 33.6∘), providing a clear but relatively long sight-line through the halo of our Galaxy. This makes the field useful for deep Galactic time-domain science with JWST as well.
This field serves as an important testbed for advancing our understanding of the time-varying universe at fainter limits than can be reached from the ground at optical and near-IR wavelengths, where transients and variability can hold cosmological significance (e.g., dark energy, cosmic star formation rate, and the evolution of supermassive blackholes). Type Ia SNe are often used as standard candles to provide critical constraints on the Hubble constant, the mass density of the universe, the cosmological constant, the deceleration parameter, and the age of the Universe (e.g., Riess et al., 1998). Core Collapse (CC) SNe are equally important, as they not only provide most of the heavy elements in our universe (e.g., Matteucci & Greggio, 1986), but also should trace the rate of instantaneous massive star formation. CC SNe rates and their connection to cosmic star formation rates remains poorly understood, as there seems to be a significant mismatch between observed CC SNe rates and what is expected (see, e.g., Cappellaro, 2014).
@@
While previous HST variability studies like Sarajedini et al. (2003) (Hubble Deep Field North) and Cohen et al. (2006) (Hubble Ultra Deep Field) achieved impressive depths (29.0 and 30.5 mag, respectively), they are limited to approximately 6.5 arcmin2. Villforth et al. (2010), on the other hand, leveraged the expansive area (500 arcmin2) of the Great Observatories Origins Deep Survey (GOODS), but to a relatively shallow depth (26.0 mag). At the same exquisite angular resolution (full width half max 009), TREASUREHUNT observations provide a unique combination of depth ( 29.5 mag) and area (88 arcmin2) for a variability analysis.
In this work, we provide the first time-domain study of the JWST NEP TDF at visible wavelengths, presenting 12 transients (supernova candidates) and 100 variable candidates (AGN candidates) identified using TREASUREHUNT data. In § 2, we briefly summarize the TREASUREHUNT HST data, and data processing specific to the detection of transients and variable objects. § 3 explains our methods for identifying transients and variability in the NEP TDF. In § 4.1 and § 4.2 we showcase our supernova candidate detections and our variable AGN candidates, respectively. All magnitudes are in AB units222Defined as mag, where the flux density is in Jy. (Oke & Gunn, 1983). Where relevant, we adopt a flat CDM cosmology with km s-1 Mpc-1, , and (Planck Collaboration et al., 2016, 2020).
@@ 2.1 HST Observations and Data Processing
Observations from HST GO 15278 were taken between 2017 October 1 and 2019 February 9, and those from GO 16252+16793 (TREASUREHUNT333We will loosely refer to the observations and data of the GO 15278 program as also part of TREASUREHUNT hereafter, as the intent and survey strategy were identical.) between 2020 September 25 and 2022 October 31. The former program targeted the central ( 5′) portion of the JWST NEP TDF, and the latter an outer annulus to 78. Together, these programs provide near-contiguous coverage with the Wide Field Camera 3 (WFC3/UVIS) in the F275W filter ( 0.272 µm) and with the Advanced Camera for Surveys (ACS/WFC) in the F435W and F606W filters ( 0.433 and 0.592 µm), with a total area of 194 arcmin2 (ACS/WFC). The full HST coverage of the JWST NEP TDF is shown in Figure 1.
As described in more detail in Jansen et al. (2024a,b; in prep.), both programs used 4-orbit CVZ and pseudo-CVZ444True CVZ orbits imply Earth Limb angles 15.5∘, but uninterrupted observations can be implemented at smaller angles down to 7.6∘ away from the Earth limb when phased to execute when that limb is dark. We term these possible but non-standard opportunities “pseudo-CVZ”. visits to reach 2- limiting depths of 28.0, 28.6, and 29.5 mag in F275W, F435W and F606W, respectively. These limiting depths were determined using the SourceExtractor (Bertin & Arnouts, 1996) MAG_AUTO aperture, within areas with single-epoch coverage (lighter grey in Figure 1), and are the adopted nominal survey depths.
@@
The background-subtracted images were corrected for geometric distortion and aligned to the Gaia/DR3 (Gaia Collaboration et al., 2016, 2023) astrometric reference grid using the DrizzlePac tasks TweakReg and AstroDrizzle (see Hoffmann et al., 2021). The AstroDrizzle software also flags cosmic-rays, a crucial step when identifying faint variability, where cosmic ray induced signals could cause spurious variability if not properly removed. We expect, and tested (as discussed in more detail in the Appendix), that all cosmic ray and other image artifacts were largely removed through the AstroDrizzle step, as each single 4-orbit visit resulted in 8–9 exposures per filter that were stacked together.
All 8–9 exposures per visit were taken at the same pointing with small 0.2″ dither offsets in RA and Dec. For each visit, we drizzle-combined the individual post-processed, rectified and aligned exposures per filter (Jansen et al. 2024a,b; in prep.). For ACS/WFC F606W and F435W, this resulted in 24 rectified and aligned science images (and 24 associated weight maps), and for WFC3/UVIS F275W in 23 images and associated weight maps, each at a pixel scale of 0030 pix-1 and centered on the nominal field center of the JWST NEP TDF. We also use the full mosaics of all drizzle-combined images per filter, generated with the same image dimensions and pixel scale, and aligned to Gaia/DR3 to within 0008 in RA and Dec.
The footprints of individual ACS/WFC (and to a lesser extent WFC3/UVIS) visits partially overlap one another, with a cumulative area of overlap of 88 arcmin2. These overlapping regions are key to searching for transients and variability within this field. Figure 1 showcases the overlapping regions in progressively darker shades of grey for areas of 2-, 3- and 4-epoch overlap. The time interval between observations in these regions of overlap range between 1 day and 4 years.
We note that while the very small number of epochs and sparse time sampling can catch transients like SNe and variable sources like AGN, they cannot provide a light curve to trace their evolution. Further epochs would provide a more robust identification of variable sources, where we may characterize the scatter in the measured brightness (e.g., Pouliasis et al., 2019; Sokolovsky et al., 2017).
@@ 2.2 JWST Observations and Data Processing
JWST/NIRCam observations within the JWST NEP TDF were obtained between 2022 August 25 and 2023 May 30 as part of JWST GTO 2738 (PIs: R. Windhorst & H. Hammel) in 7 broadband filters (F090W through F444W) and 1 mediumband filter (F410M) as part of the PEARLS program (Windhorst et al., 2023). Figure 1 shows the JWST observations outlined in orange, highlighting the 4 epochs, corresponding to the four orthogonal “spokes”. Each spoke consists of two NIRCam pointings that partially overlap in the center. Each pointing includes a total exposure time of 3.5 hours across all 8 filters, or about 3000 sec per filter. Each spoke has an area of 137, totaling 547 of JWST observations.
The data were retrieved from MAST and post-processed by the PEARLS team using their custom pipeline to mitigate noise, alleviate wisps in the NIRCam/SW filters, mask snowball artifacts, and flatten the background across read-out amplifier boundaries. The individual post-processed images where rectified and aligned to Gaia/DR3 independent of the HST images, but using nearly identical methods. Full mosaics of the field were created for each filter with a 003 platescale. The 5 sigma point-source limit is typically between 28.0 and 29.1 mag, depending on the filter. For more details on the reduction, calibration and post-processing, we refer to Windhorst et al. (2023), Robotham et al. (2023), and Jansen et al. 2024c (in prep.). In § 4.3, we will use PEARLS aperture photometry of transient and varying objects discovered in areas with both HST and JWST coverage to fit SEDs.
@@ 2.3 X-Ray Observations and Data Processing
The JWST NEP TDF has also been observed extensively in X-rays with NuSTAR, XMM-Newton, and Chandra. NuSTAR has monitored the field since Cycle 5 (PID 5192, 6218, 8180, and 9267; PI: F. Civano). For this work we use the Cycle 5+6 data and catalog published in Zhao et al. (2021a, b, 2024). The 21 NuSTAR observations took place over 28 months (between 2019 September and 2022 January), while the three XMM-Newton observations (part of the NuSTAR cycle 6 program) were taken over 15 months (between 2020 October and 2022 January). The total NuSTAR exposure from Cycle 5+6 is 1.5 Ms, reaching a flux limit of 3.310-14 erg cm-2 s-1 in the 3–24 keV band. We note that this is the deepest NuSTAR field taken thus far. The total XMM-Newton exposure time is 62 ks, reaching a limit of 8.710-16 erg cm-2 s-1 in the 0.5–2 keV band.
We also use the first 1.3 Ms of Chandra observations of the JWST NEP TDF (PID 19900666, 20900658, 21900294 and 22900038; PI: W.P. Maksym) spanning cycles 19–23. Chandra observed the field 46 times between 2018 April 12 and 2022 September 25 (a span of 4.6 years), with exposures ranging from 9 to 59.9 ks. Beginning 2021 August 29, these occurred in 90 ks epochs every 3 months, typically broken into 2 or more individual visits (OBSIDs) spanning no more than 30 days per epoch. We reprocessed and reduced the data using ciao (Fruscione et al., 2006), corrected the cross-observations using source catalogs generated by wavdetect, and astrometrically registered them to Pan-STARRS (Chambers et al., 2016). We then used merge_obs to generate combined event files and exposure maps. The faintest 3 detection in the catalog is erg s-1 cm-2 in the full 0.5–7 keV band assuming galactic absorption and a power law.
@@ 3 Methods to Identify Transients and Variability
In this section, we detail our methods for identifying transients and variability in the 55 areas of overlap (shown as darker shades of grey in Figure 1) with a combined area of 88 arcmin2.
@@ 3.2 Variability Search
Note. For each SourceExtractor parameter in the first column, the second an third columns list its value for identifying variable sources and for full aperture photometry, respectively. We used a circular aperture with a diameter of 024 diameter for variability identification, specified using the PHOT_APERTURES parameter, and the FLUX_AUTO aperture for full aperture photometry. We set the GAIN value to the median exposure time of the individual exposures combined per visit.
@@ 3.2.3 Accounting for False Detections
In addition to false detections due to Gaussian statistics, image artifacts like cosmic rays, hot pixels, bias stripping, amplifier offsets, CTE trails, dust motes, optical ghosts, and cross-talk, may cause artificial variability to be detected. Of these, cosmic rays are the brightest and most prevalent, yet will not affect our results due to robust flagging and having 8 exposures (per pointing, filter and epoch) to stack using the AstroDrizzle software. We explore this in more detail in the Appendix. The other image artifacts are largely removed by the standard HST pipeline. Therefore, we do not expect and find no evidence for image artifacts to bias our results. Nonetheless, confirmation of which individual sources genuinely vary will require follow-up observations.
@@ 4.1 Transients in the JWST NEP Time Domain Field
In all, we detected 0.14 transients per arcmin2 of overlapping area, or 491 deg-2. Since two epochs are needed to detect a transient, the areal density per epoch is therefore 0.07 transients per arcmin2 (245 deg-2), or 0.77 per ACS/WFC footprint per epoch. For comparison, Dahlen et al. (2012) searched for SNe in the GOODS fields (using the ACS/F850LP filter), and identified 118 SNe in a total area of 1 deg2 to 26 mag. If we only consider the 4 transients with 26.0 mag, we find an areal density of 164 deg-2, consistent with Dahlen et al. (2012) when factoring in differences in detection filter, SNe colors and redshift distribution, and small number statistics. We also note that it is likely that not all our transients are SNe.